OpenStudy (anonymous):

A cylindrical tank has a radius of 23 m and a height of 52 m. What is its lateral area? Use 3.14 for .

OpenStudy (anonymous):

do you know what the lateral area of a cylinder is?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

 the lateral area is the curved side of the cylinder

OpenStudy (anonymous):

how do i do the problem?

OpenStudy (anonymous):

basically lateral area can be solved by multiplying the height of the cylinder by the circumference of the circle

OpenStudy (anonymous):

so 23 +23 = 46 so its 46 x 52?

OpenStudy (anonymous):

circumference= 2*pi*r

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

23+23 = diameter circumference is also d*pi

OpenStudy (anonymous):

2x 3.14 x 23?

OpenStudy (anonymous):

then multiply that by the height to get your answer

OpenStudy (anonymous):

7510?

OpenStudy (anonymous):

no idea, i didnt calculate it out

OpenStudy (anonymous):

thats an answer so its possible! thank you

OpenStudy (anonymous):

i got 10/10 thank you ")